# On the Cutoff Scale Identification of FLRW Cosmology in Asymptotically Safe Gravity

### v1

@@ III.3 The identification by the Hubble parameter with quantum correction
$k(H)=\zeta G_{0}^{(\beta-1)/2}H^{\beta},$ (31)
where $\zeta$ is a dimensionless constant. To ensure that $\zeta$ in the cutoff scale is dimensionless, unlike in [24], we introduce $G_{0}$ within the cutoff scale. Three specific values of $\beta=1/4,3/4,1$ were considered in [24], and we study which of these values are viable. By substituting this cutoff choice into (3) with $\Lambda_{0}=0$, we obtain the leading correction to the cosmological constant as
@@
where $\gamma=\zeta^{4}\nu G_{0}^{2\beta-1}/3$. Putting (32) in (15), the Einstein equation (15) for $H(t)$ becomes
@@
$-\frac{H^{2-4\beta}}{H\left(\gamma-H^{2-4\beta}\right)}dH=-\alpha\frac{da}{a}~{}.$ (34)
@@
For the sake of completeness, let us examine the cosmic evolution when the exponent of the momentum cutoff scale is exactly $\beta=1/2$. This gives $k=\zeta G_{0}^{-1/4}H^{1/2}=\zeta\sqrt{m_{\mathrm{Pl}}}H^{1/2}$ and the leading order cosmological constant becomes $\Lambda=\zeta^{4}\nu H^{2}$. The Friedmann equation then takes the form
$\frac{dH}{dt}=-\alpha\left(1-\frac{\zeta^{4}\nu}{3}\right)H^{2}~{}.$ (42)
@@
$a(t)=a_{0}(t-c)^{\frac{1}{\alpha(1-\zeta^{4}\nu/3)}}~{}.$ (44)
We observe that the solution maintains a power-law behaviour, similar to the classical case, unlike the solutions for $\beta<1/2$. Although the cutoff scale $k$ is large, scaling with the Planck mass, the running cosmological constant does not depend on the Planck mass and remains small. Thus, $\beta\geq 1/2$ is a viable cutoff choice and $\beta=1/2$ represents the critical value, below which $\beta$ are not good choices for late time.
@@ IV.1 Classical solution
To calculate the classical solution for the Hubble parameter $H_{\mathrm{cl}}(t)$ in a universe dominated by a cosmological constant, the classical Einstein equation (15) with $\Lambda=\Lambda_{0}>0$ yields:
@@ IV.2 The identification using the cosmic time
which was successful for $\Lambda_{0}=0$. By inserting this cutoff scale in Eq. (3), we can express the running gravitational coupling and cosmological coupling as functions of time as
@@
By setting the quantum parameters to zero, we should recover the leading-order classical term of the Hubble parameter. However, the classical solution shows the exponential expansion, which can never reproduced from the power expansion in $t$. Another important observation is that, unlike the $\Lambda_{0}=0$ case, this solution lacks an undetermined constant. This absence is unusual because the Einstein equation is a first-order differential equation for $H(t)$, which typically permits a constant of integration. All of these suggest that the cutoff identification $k=\xi/t$ is not a physically viable choice.
@@ IV.3 The identification using the scale factor
where we have used the cosmological constant to introduce dimension, and we use this for identification
@@
With the choice of cutoff (55), the scale-dependent cosmological constant (3) becomes
@@ V Summary and Conclusions
In this work, we have studied a spatially flat FLRW universe in late time, taking quantum gravitational effects into account, within the framework of asymptotically safe gravity. The FRG flow of the effective average action in asymptotically safe gravity has resulted in the running of the gravitational constant and the cosmological constant under the Einstein-Hilbert truncation. We have analyzed the cosmic evolution of the FLRW cosmology at late times by improving the Einstein equations to include the scale dependence of Newton’s constant and the cosmological constant. Specifically, we have considered the conservation of the entire right-hand side of the Einstein equations, which has led to a modified continuity equation. The main focus of this work has been to identify a viable cutoff scale, $k$, that can yield a consistent FLRW cosmology.
In this context, we have examined cosmological solutions for two branches of FRG trajectories. The first branch has involved a power series expansion of the dimensionful cosmological constant, $\Lambda(t)$, beginning with a $k^{4}$ term, with $\Lambda_{0}=0$. This trajectory has approached the Gaussian fixed point as $k\rightarrow 0$ and has been referred to in the literature as the “separatrix” [8, 6]. The second branch, which may be realized in nature, starts with a $k^{2}$ term in the expansion of the cosmological constant and is valid up to a certain non-zero momentum scale, where the $\beta$-function becomes singular as the dimensionless cosmological constant $\lambda(k)$ comes to a “singular line” near $1/2$.
@@
Considering another possibility that the cutoff scale $k$ may have a simple functional relationship with the quantum Hubble parameter rather than with the classical Hubble parameter, we examined the cutoff choice $k(H)=\zeta G_{0}^{(\beta-1)/2}H^{\beta}$, where $\beta$ is a constant. In earlier work [24], three specific values of $\beta$ namely, $\beta=1/4,3/4$ and 1 were analyzed, and the corresponding entropy generation due to quantum effects was investigated. It was observed that entropy generation diverges at late times for $\beta=1/4$, indicating that this cutoff choice is not suitable for studying cosmology. On the other hand, for $\beta=3/4$ and $\beta=1$, entropy generation converges to a constant and thus these viable cutoff choices are viable. This suggests the existence of a critical value of $\beta$ that determines the appropriateness of a cutoff choice. By studying the asymptotic behaviour of the Hubble parameter at large times, we have demonstrated that the critical value is $\beta=1/2$. This value marks a transition in the scale factor’s behaviour from power-law growth ($\beta\geq 1/2$) to exponential growth ($\beta<1/2$). From the exponential behaviour of the scale factor for $\beta<1/2$, we infer that such cutoff choices are not viable, as the large momentum scale $k$ introduces significant perturbations that entirely disrupt the classical power-law behaviour for the $\Lambda_{0}=0$ trajectory. On the other hand, for $\beta\geq 1/2$, the behaviour follows a power law similar to the classical case. Here the perturbations remain small enough to preserve this expected behaviour with minor quantum corrections. Therefore, the cutoff identification is valid for $\beta\geq 1/2$ making it suitable for cosmological studies.
@@ Appendix A Perturbative Solution of Running Couplings
$k\partial_{k}\Lambda(k)$ $=$ $\eta_{N}\Lambda+\frac{k^{4}G(k)}{2\pi}\left[10\Phi^{1}_{2}(-2\lambda)-8\Phi^{1}_{2}(0)-5\eta_{N}\tilde{\Phi}^{1}_{2}(-2\lambda)\right],$ (67)
@@
$\eta_{N}=\frac{gB_{1}(\lambda)}{1-gB_{2}(\lambda)},$ (68)
in which the dimensionless running couplings are defined as $g(k)=G(k)k^{2},\lambda(k)=\Lambda(k)/k^{2}$ and $B_{1}(\lambda),B_{2}(\lambda)$ are abbreviations of
@@
The threshold functions that appeared here are defined as
[…]
which depend on the choice the cutoff function $R^{(0)}(z)$.
@@ A.1 Exponential Cutoff
Case I: $\Lambda_{0}=\Lambda_{2}=0$ and $\Lambda_{i}=G_{i}=0$ for odd $n$
@@
${\tilde{\Phi}}^{1}_{1}(0)=1,\qquad{\tilde{\Phi}}^{1}_{2}(0)=1,\qquad{\tilde{\Phi}}^{2}_{1}(0)=\ln 2,\qquad{\tilde{\Phi}}^{2}_{2}(0)=\frac{1}{2}.$ (80)
The solutions agree with the results in [7] as, in the notation of (3): $G_{2}=-\omega G_{0}^{2},G_{4}=\omega_{1}G_{0}^{3},\Lambda_{4}=\nu G_{0},\Lambda_{6}=\nu_{1}G_{0}^{2}$,
@@
$+$ $\frac{(-1)^{p}}{2^{p+1}}\left(\frac{p(p+1)(p+2)\Lambda_{1}^{3}+6p\Lambda_{0}\Lambda_{3}-6p(p+1)\Lambda_{0}\Lambda_{1}\Lambda_{2}}{3\Lambda_{0}^{p+3}}\Phi^{0}_{n}(0)+\frac{p(p+1)\Lambda_{1}}{\Lambda_{0}^{p+2}}\Phi^{-1}_{n}(0)\right)k^{2p+3}+\cdots,$
@@
$\Lambda_{2}=-\frac{2\Phi^{1}_{1}(0)+3\Phi^{2}_{2}(0)}{3\pi}G_{0}\Lambda_{0},\qquad\Lambda_{4}=\frac{\left[2\Phi^{1}_{1}(0)+3\Phi^{2}_{2}(0)\right]^{2}}{9\pi^{2}}G_{0}^{2}\Lambda_{0}-\frac{5\Phi^{0}_{1}(0)+24\Phi^{1}_{2}(0)}{24\pi}G_{0}.$ (84)
With the specific values (A.1) we have
$G(k)$ $=$ $G_{0}\left[1-\frac{9+\pi^{2}}{9\pi}G_{0}k^{2}+\left[\left(\frac{9+\pi^{2}}{9\pi}\right)^{2}-\frac{5\pi}{72}\frac{1}{G_{0}\Lambda_{0}}\right]G_{0}^{2}k^{4}+\mathcal{O}(k^{6})\right],$ (85)
$\Lambda(k)$ $=$ $\Lambda_{0}\left[1-\frac{9+\pi^{2}}{9\pi}G_{0}k^{2}+\left(\frac{9+\pi^{2}}{9\pi}\right)^{2}G_{0}^{2}k^{4}+\mathcal{O}(k^{6})\right]+G_{0}k^{4}\left[-\frac{5\pi^{2}+144\zeta(3)}{72\pi}+\mathcal{O}(k^{2})\right].$ (86)
This result gives the coefficients in Eq. (3) for the exponential cutoff.
@@ A.2 Optimized Cutoff
Note that these two cases are “disconnected”, i.e. the coefficients in the case $\Lambda_{0}=0$ are not the $\Lambda_{0}\to 0$ limit of the related coefficients in the case $\Lambda_{0}\neq 0$.

### v2

@@ III.3 The identification by the Hubble parameter with quantum correction
$k(H)=\xi G_{0}^{(\beta-1)/2}H^{\beta},$ (31)
where $\xi$ is a dimensionless constant. To ensure that $\xi$ in the cutoff scale is dimensionless, unlike in [25], we introduce $G_{0}$ within the cutoff scale. Three specific values of $\beta=1/4,3/4,1$ were considered in [25], and we study which of these values are viable. By substituting this cutoff choice into (3) with $\Lambda_{0}=0$, we obtain the leading correction to the cosmological parameter as
@@
where $\gamma=\xi^{4}\nu G_{0}^{2\beta-1}/3$. Putting (32) in (15), the Einstein equation (15) for $H(t)$ becomes
@@
$\frac{H^{2-4\beta}}{H\left(\gamma-H^{2-4\beta}\right)}dH=-\alpha\frac{da}{a}~.$ (34)
@@
For the sake of completeness, let us examine the cosmic evolution when the exponent of the momentum cutoff scale is exactly $\beta=1/2$. This gives $k=\xi G_{0}^{-1/4}H^{1/2}=\xi\sqrt{m_{\mathrm{Pl}}}H^{1/2}$ and the leading order cosmological parameter becomes $\Lambda=\xi^{4}\nu H^{2}$. The Friedmann equation then takes the form
$\frac{dH}{dt}=-\alpha\left(1-\frac{\xi^{4}\nu}{3}\right)H^{2}~.$ (42)
@@
$a(t)=a_{0}(t-c)^{\frac{1}{\alpha(1-\xi^{4}\nu/3)}}~.$ (44)
We observe that the solution maintains a power-law behaviour, similar to the classical case, unlike the solutions for $\beta<1/2$. Although the cutoff scale $k$ is large, scaling with the Planck mass, the running cosmological parameter does not depend on the Planck mass and remains small. Thus, $\beta\geq 1/2$ is a viable cutoff choice and $\beta=1/2$ represents the critical value, below which $\beta$ are not good choices for late time.
@@ IV.1 Classical solution
To calculate the classical solution for the Hubble parameter $H_{\mathrm{cl}}(t)$ in a universe dominated by a cosmological parameter, the classical Einstein equation (15) with $\Lambda=\Lambda_{0}>0$ yields:
@@ IV.2 The identification using the cosmic time
which was successful for $\Lambda_{0}=0$. By inserting this cutoff scale in Eq. (3), we can express the running Newton coupling and cosmological parameter as functions of time as
@@
By setting the quantum parameters to zero, we should recover the leading-order classical term of the Hubble parameter. However, the classical solution expands exponentially which can never be reproduced from the power expansion in $t$. Another important observation is that, unlike the $\Lambda_{0}=0$ case, this solution lacks an undetermined constant. This absence is unusual because the Einstein equation is a first-order differential equation for $H(t)$, which typically permits a constant of integration. All of these suggest that the cutoff identification $k=\xi/t$ is not a physically viable choice.
@@ IV.3 The identification using the scale factor
where we have used the cosmological parameter to introduce dimension, and we use this for identification
@@
With the choice of cutoff (55), the scale-dependent cosmological parameter (3) becomes
@@ V Summary and Conclusions
In this work, we have studied a spatially flat FLRW universe in late time, taking quantum gravitational effects into account, within the framework of asymptotically safe gravity. The FRG flow of the effective average action in asymptotically safe gravity has resulted in the running of the Newton coupling and the cosmological parameter under the Einstein-Hilbert truncation. We have analyzed the cosmic evolution of the FLRW cosmology at late times by improving the Einstein equations to include the scale dependence of Newton coupling and the cosmological parameter. Specifically, we have considered the conservation of the entire right-hand side of the Einstein equations, which has led to a modified continuity equation. The main focus of this work has been to identify a viable cutoff scale, $k$, that can yield a consistent FLRW cosmology.
In this context, we have examined cosmological solutions for two branches of FRG trajectories. The first branch has involved a power series expansion of the dimensionful cosmological parameter, $\Lambda(t)$, beginning with a $k^{4}$ term, with $\Lambda_{0}=0$. This trajectory has approached the Gaussian fixed point as $k\rightarrow 0$ and has been referred to in the literature as the “separatrix” [9, 7]. The second branch, which may be realized in nature, starts with a $k^{2}$ term in the expansion of the cosmological parameter and is valid up to a certain non-zero momentum scale, where the $\beta$-function becomes singular as the dimensionless cosmological parameter $\lambda\equiv\Lambda(k)/k^{2}$ comes to a “singular line” near $1/2$. (The beta function becomes singular as can be seen from Eq. (90). Similar singularity exists for other choices of the regulators. The RG flow near this singularity is discussed in detail in Ref. [10].)
@@
Considering another possibility that the cutoff scale $k$ may have a simple functional relationship with the quantum Hubble parameter rather than with the classical Hubble parameter, we examined the cutoff choice $k(H)=\xi G_{0}^{(\beta-1)/2}H^{\beta}$, where $\beta$ is a constant. In earlier work [25], three specific values of $\beta$ namely, $\beta=1/4,3/4$ and 1 were analyzed, and the corresponding entropy generation due to quantum effects was investigated. It was observed that entropy generation diverges at late times for $\beta=1/4$, indicating that this cutoff choice is not suitable for studying cosmology. On the other hand, for $\beta=3/4$ and $\beta=1$, entropy generation converges to a constant and thus these cutoff choices are viable. This suggests the existence of a critical value of $\beta$ that determines the appropriateness of a cutoff choice. By studying the asymptotic behaviour of the Hubble parameter at large times, we have demonstrated that the critical value is $\beta=1/2$. This value marks a transition in the scale factor’s behaviour from power-law growth ($\beta\geq 1/2$) to exponential growth ($\beta<1/2$). From the exponential behaviour of the scale factor for $\beta<1/2$, we infer that such cutoff choices are not viable, as the large momentum scale $k$ introduces significant perturbations that entirely disrupt the classical power-law behaviour for the $\Lambda_{0}=0$ trajectory. On the other hand, for $\beta\geq 1/2$, the behaviour follows a power law similar to the classical case. Here the perturbations remain small enough to preserve this expected behaviour with minor quantum corrections. Therefore, the cutoff identification is valid for $\beta\geq 1/2$ making it suitable for cosmological studies.
@@ Appendix A Perturbative Solution of Running Couplings
$k\partial_{k}\Lambda(k)$ $=$ $\eta_{N}\Lambda(k)+\frac{k^{4}G(k)}{2\pi}\left[10\Phi^{1}_{2}(-2\lambda)-8\Phi^{1}_{2}(0)-5\eta_{N}\tilde{\Phi}^{1}_{2}(-2\lambda)\right],$ (67)
@@
$\eta_{N}=\frac{k^{2}G(k)\,B_{1}(\lambda)}{1-k^{2}G(k)\,B_{2}(\lambda)},$ (68)
in which the dimensionless coupling is defined as $\lambda=\Lambda(k)/k^{2}$, and $B_{1}(\lambda)$ and $B_{2}(\lambda)$ are abbreviations of
@@
The threshold functions that appear here are defined as
$\Phi^{p}_{n}(w)=\frac{1}{\Gamma(n)}\int_{0}^{\infty}dzz^{n-1}\frac{R^{(0)}(z)-z\partial_{z}R^{(0)}(z)}{\left[z+R^{(0)}(z)+w\right]^{p}},\qquad{\tilde{\Phi}}^{p}_{n}(w)=\frac{1}{\Gamma(n)}\int_{0}^{\infty}dzz^{n-1}\frac{R^{(0)}(z)}{\left[z+R^{(0)}(z)+w\right]^{p}},$ (70)
which depend on the choice the cutoff function $R^{(0)}(z)$. Throughout this paper, $\Gamma(n)$ is the gamma function.
@@ A.1 Exponential Cutoff
Case I: $\Lambda_{0}=0$
@@
${\tilde{\Phi}}^{1}_{1}(0)=1,\qquad{\tilde{\Phi}}^{1}_{2}(0)=1,\qquad{\tilde{\Phi}}^{2}_{1}(0)=\ln 2,\qquad{\tilde{\Phi}}^{2}_{2}(0)=\frac{1}{2},$ (80)
where $\zeta(3)$ is the value of the zeta function $\zeta(n)$ at $n=3$. The solutions agree with the results in [8] as, in the notation of (3): we find $\Lambda_{2}=0$, $\Lambda_{n}=G_{n}=0$ for odd $n$ and $G_{2}=-\omega G_{0}^{2},G_{4}=\omega_{1}G_{0}^{3},\Lambda_{4}=\nu G_{0},\Lambda_{6}=\nu_{1}G_{0}^{2}$,
@@
$+$ $\frac{(-1)^{p}}{2^{p+1}}\left(\frac{p(p+1)(p+2)\Lambda_{1}^{3}+6p\Lambda_{0}^{2}\Lambda_{3}-6p(p+1)\Lambda_{0}\Lambda_{1}\Lambda_{2}}{3\Lambda_{0}^{p+3}}\Phi^{0}_{n}(0)+\frac{p(p+1)\Lambda_{1}}{\Lambda_{0}^{p+2}}\Phi^{-1}_{n}(0)\right)k^{2p+3}+\cdots,$
@@
$\Lambda_{2}=-\frac{2\Phi^{1}_{1}(0)+3\Phi^{2}_{2}(0)}{3\pi}G_{0}\Lambda_{0},\qquad\Lambda_{4}=\frac{\left[2\Phi^{1}_{1}(0)+3\Phi^{2}_{2}(0)\right]^{2}}{9\pi^{2}}G_{0}^{2}\Lambda_{0}-\frac{5\Phi^{0}_{1}(0)+24\Phi^{1}_{2}(0)}{24\pi}G_{0},$ (84)
$\Lambda_{6}=-\frac{\left[2\Phi^{1}_{1}(0)+3\Phi^{2}_{2}(0)\right]^{3}G_{0}^{3}\Lambda_{0}}{27\pi^{3}}+\frac{\left[2\Phi^{1}_{1}(0)+3\Phi^{2}_{2}(0)\right]\left[10\Phi^{0}_{1}(0)+36\Phi^{1}_{2}(0)-5\tilde{\Phi}^{0}_{1}(0)\right]G_{0}^{2}}{108\pi^{2}}-\frac{\left[48\Phi^{0}_{2}(0)+5\Phi^{-1}_{1}(0)\right]G_{0}}{72\pi\Lambda_{0}}.$
With the specific values (A.1) and
$\Phi^{-1}_{1}(0)=\frac{\pi^{2}}{2}+3\zeta(3),\qquad\Phi^{0}_{2}(0)=6\zeta(3),\qquad\tilde{\Phi}^{0}_{1}(0)=\frac{\pi^{2}}{6},$ (85)
we have
$G(k)$ $=$ $G_{0}\left[1-\frac{9+\pi^{2}}{9\pi}G_{0}k^{2}+\left[\left(\frac{9+\pi^{2}}{9\pi}\right)^{2}-\frac{5\pi}{72}\frac{1}{G_{0}\Lambda_{0}}\right]G_{0}^{2}k^{4}+\mathcal{O}(G_{0}^{3}k^{6})\right],$ (86)
$\Lambda(k)$ $=$ $\Lambda_{0}\left[1-\frac{9+\pi^{2}}{9\pi}G_{0}k^{2}+\left(\frac{9+\pi^{2}}{9\pi}\right)^{2}G_{0}^{2}k^{4}-\left(\left(\frac{9+\pi^{2}}{9\pi}\right)^{3}+\frac{5\pi^{2}+606\zeta(3)}{144\pi\Lambda_{0}^{2}G_{0}^{2}}\right)G_{0}^{3}k^{6}+\mathcal{O}(G_{0}^{4}k^{8})\right]$ (87)
$+G_{0}k^{4}\left[-\frac{5\pi^{2}+144\zeta(3)}{72\pi}+\frac{(9+\pi^{2})(5\pi^{2}+144\zeta(3))}{648\pi^{2}}G_{0}k^{2}+\mathcal{O}(G_{0}^{2}k^{4})\right].$
This result gives the coefficients in Eq. (3) for the exponential cutoff. Note that these two cases are “disconnected”, i.e. the coefficients in the case $\Lambda_{0}=0$ are not the $\Lambda_{0}\to 0$ limit of the related coefficients in the case $\Lambda_{0}\neq 0$.
@@ A.2 Optimized Cutoff
Again these two cases are “disconnected” as in the exponential cutoff.